Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Sensing and control implementation)
ruggi, vardaro - 17:55 Monday 24 January 2022 (54585) Print this report
Extraction of a signal for DARM Optical Spring pole frequency

In the shift of last Monday a channel for the monitor of the optical spring pole is added. The channel compute the transfer function between LSC_DARM and SNEB_B7_DC at the frequency of the DARM UGF Line i.e. 74.4 Hz. The name of the channel is LSC_OS_LF_mag.

After that this channel was added in the data flow we observed the following things:

  • The LSC_OS_LF_mag as expected senses different kind of adjustment of the ITF working point (see fig 1)
  • During the scan of SR_TX angular alignment this signal remained constant as expected (see fig 2.) (see elog 54460) 
  • During the scan of SR_TY this channels changed as expected  (see fig 3.) (see elog 54460)
  • At the end of the working point tuning we did a longitudinal scan of the SR longitudinal loop in order to have the oprtical spring pole around zero and we profit of this scan for the calibration of the channel

Figure 4 shows:

  • the LSC_OS_LF_mag channel during the SRCL_SET scan that decreased from 40 to 6 (up left polt),
  • the LSC_OS_LF_mag channel afrer the calibration (up right plot)
  • the LSC_OS_LF_phi is the phase monitor of the channel but it is too noisy (bottom left plot)
  • the SRCL_SET channel (bottom right plot)

With each points of SRCL set we measured the transfer function of DARM (fig 5) and we fitted the OS pole frequency:

  • SRCL SET = 0.0       pole freq =   14.6 Hz
  • SRCL SET = 1.0       pole freq =  13.7 Hz
  • SRCL SET = 2.0       pole freq =  12.5 Hz
  • SRCL SET = 4.0       pole freq =  10.1 Hz
  • SRCL SET = 6.0       pole freq =   7.8  Hz
  • SRCL SET = 6.5       pole freq =    6.3 Hz

we had some difficulties to fit the last two poles. We checked also the transfer function between LSC_DARM_ERR and SNEB_B7_DC and we saw that it scales with the set point of SRCL and that in the last two points the coherence is lower than with the other poitns (see fig 6).

For each SRCL_SET we averaged the LSC_OS_LF_mag and LSC_OS_LF_phi channels and we saw (figure 7)

  • The LSC_OS_LF_phi behaves randomly with the SRCL set thus it is not trustable (secnod row)
  • The LSC_OS_LF_mag scales linearly with the SRCL_SET except the first point (first row)
  • The sqrt(LSC_OS_LF_mag) scale linearly with the frequency of the Optical spring pole with the following law sqrt(LSC_OS_LF_mag) = m*freq_pole with m=0.40+/-0.03

We did the same exercise using the SNEB_B7_DC demodulated at the frequency of the DARM line i.e. 74.4 Hz. The results are shown in fig 8 and 9. The name of this channel is LSC_NArm_OS_mag. The linear coefficent is 0.00120+/-0.00005

We did a comparison plot (fig 10) between the LSC_NArm_OS_mag and LSC_OS_LF_mag converted in Hz and we realized that the signal extracted without passing to the TF with DARM is cleaner with higher OS pole frequency.



We added in the data the channel LSC_OS_LF_mag_Hz that is the LSC_OS_LF_mag signal calinrated in Hz (frequency of the pole of the optical spring).  This signal has 2 issues:

  1. it is signless because it is the magnitude of the DARM line in SNEB_B7_DC signal and the phi channel is not trustable
  2. it is can not sense optical springs below 6 Hz

We are working on another way to extract this signal. Further details will follow

Images attached to this report
Search Help